# LOOPerSet: A Large-Scale Dataset for Data-Driven Polyhedral Compiler Optimization

### v1

@@ 1 Introduction
The polyhedral model [8, 15] offers a powerful framework for expressing and applying complex loop transformations, which are essential for optimizing scientific and high-performance applications. The central challenge, however, lies in navigating the vast search space of legal transformation sequences to find one that delivers optimal performance on a given hardware target. For years, this complex search has been effectively guided by established analytical cost models and sophisticated heuristics [4, 5, 9, 12, 14]. These methods are designed to be general and tractable, typically optimizing for key performance proxies like improving data locality and enabling parallelism.
A different and complementary approach is to replace these analytical models with empirical ones built using machine learning. By training on large amounts of performance data, a learned model can develop a more nuanced understanding of the true cost and benefit of transformations on real hardware. It can learn to capture the subtle, non-linear interactions between optimizations and the underlying system that are difficult to model analytically. This data-driven approach, validated by prior work like the Tiramisu autoscheduler [2] and others [6, 18, 1, 11], has shown great potential for discovering highly effective optimization strategies.
@@
The dataset was originally created to develop and train the LOOPer autoscheduler [10]. While that paper details the architecture and evaluation of the LOOPer system, this paper serves as the reference for the dataset itself. We begin by providing a high-level overview of the dataset in Section 2. We then detail our relevance-guided data generation process in Section 3 and present a rigorous statistical characterization and quantitative diversity analysis in Section 4 to validate the dataset’s quality. Finally, Section 6 provides instructions on the dataset’s format, licensing, and access to enable its immediate use by the research community.
@@ 2 Dataset Overview
The full dataset is publicly available and free to use for academic and commercial research. It is accessible on Hugging Face [^1] https://huggingface.co/datasets/Mascinissa/LOOPerSet along with tools and examples for parsing the data. Researchers can use either the full 28-million-point dataset or a smaller 10-million-point version that was used for the experiments in the LOOPer paper [10].
@@ 3.2 Transformation Space Sampling: Relevance-Guided Exploration
To achieve this, we employed a relevance-guided sampling strategy. This approach uses the execution-guided search mechanism of the LOOPer autoscheduler [10] to explore the transformation space for each synthetic program. This mechanism performs a beam search to explore promising sequences of key polyhedral optimizations, including loop fusion, skewing, interchange, reversal, tiling, parallelization, and unrolling.
@@
Every transformation sequence considered during this process was verified for correctness using standard polyhedral dependency analysis [16, 7] to ensure that only legal (i.e., semantics-preserving) schedules were included in the final dataset.
@@ 3.3 Performance Label Generation
The final stage of the pipeline generates a ground-truth performance label for each (program, schedule) pair. Using the Tiramisu compiler framework [3], we applied the specified transformations to generate an executable file for each pair. The resulting object file was then linked with a measurement wrapper and run on our target hardware, a dual-socket system with Intel Xeon E5-2695 v2 processors. To ensure our measurements were stable and robust against system noise, we executed each program version up to 30 times and recorded the full list of execution times. This entire process was a significant computational undertaking, requiring tens of thousands CPU-hours of machine time to generate the full dataset.
While the performance labels in this dataset are specific to this CPU architecture, the work on the LOOPer system [10] has shown that the underlying modeling techniques can be adapted and ported to other hardware targets. The dataset is also well-suited for pre-training. A model can be trained on our large dataset to learn the general relationships between program structures, transformations, and performance. This pre-trained model can serve as a starting point that can be adapted to a new target architecture through transfer learning or fine-tuning, using only a much smaller, more targeted set of measurements from the new hardware. This approach helps mitigate the "cold start" problem that complicates the development of learned optimizers for new architectures.
@@ 4.2 Quantitative Diversity Analysis
A key question for any synthetic dataset is whether it truly explores new ground or simply re-creates existing examples. To answer this, we performed a quantitative analysis to assess the diversity of our programs and to formally check for any overlap with the PolyBench suite [13].
@@ 4.2.1 Methodology: Measuring Program Similarity
To quantitatively analyze our dataset’s diversity and address the possibility of benchmark replication, we needed a formal way to measure the similarity between any two programs. For this, we chose the normalized Tree Edit Distance (nTED)[17] as our primary metric .
@@ 6.1 Availability and Access
  • A smaller 10-million-point version, which is pre-split into a training set and a validation set. This split is identical to the one used to train and evaluate the LOOPer cost model [10], enabling direct replication of its main results.

### v2

@@ 1 Introduction
The polyhedral model [Feautrier2011, thangamani_survey_2024] offers a powerful framework for expressing and applying complex loop transformations, which are essential for optimizing scientific and high-performance applications. The central challenge, however, lies in navigating the vast search space of legal transformation sequences to find one that delivers optimal performance on a given hardware target. For years, this complex search has been effectively guided by established analytical cost models and sophisticated heuristics [10.1145/2896389, bondhugula_practical_2008, grosser_polly_2012, pop_graphite_2006, pouchet_polyopt_2012]. These methods are designed to be general and tractable, typically optimizing for key performance proxies like improving data locality and enabling parallelism.
A different and complementary approach is to replace these analytical models with empirical ones built using machine learning. By training on large amounts of performance data, a learned model can develop a more nuanced understanding of the true cost and benefit of transformations on real hardware. It can learn to capture the subtle, non-linear interactions between optimizations and the underlying system that are difficult to model analytically. This data-driven approach, validated by prior work like the Tiramisu autoscheduler [baghdadi_deep_2021] and others [chen_learning_2018, zheng_ansor_2020, adams_learning_2019, phothilimthana_tpugraphs_2023], has shown great potential for discovering highly effective optimization strategies.
@@
The dataset was originally created to develop and train the LOOPer autoscheduler [looper_pact25]. While that paper details the architecture and evaluation of the LOOPer system, this paper serves as the reference for the dataset itself. We begin by providing a high-level overview of the dataset in Section 2. We then detail our relevance-guided data generation process in Section 3 and present a rigorous statistical characterization and quantitative diversity analysis in Section 4 to validate the dataset’s quality. Finally, Section 6 provides instructions on the dataset’s format, licensing, and access to enable its immediate use by the research community.
@@ 2 Dataset Overview
The full dataset is publicly available and free to use for academic and commercial research. It is accessible on Hugging Face [^1] https://huggingface.co/datasets/Mascinissa/LOOPerSet along with tools and examples for parsing the data. Researchers can use either the full 28-million-point dataset or a smaller 10-million-point version that was used for the experiments in the LOOPer paper [looper_pact25].
@@ 3.2 Transformation Space Sampling: Relevance-Guided Exploration
To achieve this, we employed a relevance-guided sampling strategy. This approach uses the execution-guided search mechanism of the LOOPer autoscheduler [looper_pact25] to explore the transformation space for each synthetic program. This mechanism performs a beam search to explore promising sequences of key polyhedral optimizations, including loop fusion, skewing, interchange, reversal, tiling, parallelization, and unrolling.
@@
Every transformation sequence considered during this process was verified for correctness using standard polyhedral dependency analysis [vasilache_violated_2006, feautrier_array_1988] to ensure that only legal (i.e., semantics-preserving) schedules were included in the final dataset.
@@ 3.3 Performance Label Generation
The final stage of the pipeline generates a ground-truth performance label for each (program, schedule) pair. Using the Tiramisu compiler framework [baghdadi_tiramisu_2019], we applied the specified transformations to generate an executable file for each pair. The resulting object file was then linked with a measurement wrapper and run on our target hardware, a dual-socket system with Intel Xeon E5-2695 v2 processors. To ensure our measurements were stable and robust against system noise, we executed each program version up to 30 times and recorded the full list of execution times. This entire process was a significant computational undertaking, requiring tens of thousands CPU-hours of machine time to generate the full dataset.
While the performance labels in this dataset are specific to this CPU architecture, the work on the LOOPer system [looper_pact25] has shown that the underlying modeling techniques can be adapted and ported to other hardware targets. The dataset is also well-suited for pre-training. A model can be trained on our large dataset to learn the general relationships between program structures, transformations, and performance. This pre-trained model can serve as a starting point that can be adapted to a new target architecture through transfer learning or fine-tuning, using only a much smaller, more targeted set of measurements from the new hardware. This approach helps mitigate the "cold start" problem that complicates the development of learned optimizers for new architectures.
@@ 4.2 Quantitative Diversity Analysis
A key question for any synthetic dataset is whether it truly explores new ground or simply re-creates existing examples. To answer this, we performed a quantitative analysis to assess the diversity of our programs and to formally check for any overlap with the PolyBench suite [pouchet_polybench_2010].
@@ 4.2.1 Methodology: Measuring Program Similarity
To quantitatively analyze our dataset’s diversity and address the possibility of benchmark replication, we needed a formal way to measure the similarity between any two programs. For this, we chose the normalized Tree Edit Distance (nTED)[doi:10.1137/0218082] as our primary metric .
@@ 6.1 Availability and Access
  • A smaller 10-million-point version, which is pre-split into a training set and a validation set. This split is identical to the one used to train and evaluate the LOOPer cost model [looper_pact25], enabling direct replication of its main results.
@@ 7 Conclusion
Acknowledgments
This research was partly supported by the Center for Artificial Intelligence and Robotics (CAIR) at New York University Abu Dhabi, funded by Tamkeen under the NYUAD Research Institute Award CG010. The authors also wish to thank the Commit research group, led by Professor Saman Amarasinghe in the Computer Science and Artificial Intelligence Laboratory (CSAIL) at MIT, for providing access to their computing cluster for the data generation process.